OpenStudy (anonymous):

how do you solve 2l3x-1l=22 (Solve for x)

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

divide by 2 and then solve two separate equations

OpenStudy (anonymous):

\[2|x=3x-1|=22\] \[|3x-1|=11\] \[3x-1=11\] or \[3x-1=-11\]

OpenStudy (anonymous):

typo on the first line should just say \[2|3x-1|=22\]

OpenStudy (anonymous):

both equation easy to solve \[3x-1=11\] \[3x=12\] \[x=4\] or \[3x-1=-11\] \[3x=-10\] \[x=-\frac{10}{3}\] and those are the two solution,

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

thanks man.

OpenStudy (anonymous):

welcome!

OpenStudy (anonymous):

\[2|3x-1|=22\] then you subtract the 2 from the 22 after this you now have \[|3x-1|=20\] now you take this and make two solutions one being 3x-1=20 and 3x-1=-20 you solv for x and your two solutions will be (7,-19/3)